OpenStudy (anonymous):

Solve: 5^4-x = 1/5 How would I go about making the bases equal, I can't find any powers that I think would work...

jimthompson5910 (jim_thompson5910):

hint: 1/5 = 5^(-1)

OpenStudy (lgbasallote):

5^(4-x) = 5^ -1 so 4-x = -1 so x = 5

OpenStudy (anonymous):

Ok, so can I add an exponent to only 1 side?

jimthompson5910 (jim_thompson5910):

not sure what you mean by that...

OpenStudy (anonymous):

I got it guys, thanks alot. I've been out of math for 10 years and tryin to get back what I knew back then...its tough lol